What is the probability that the sum is less than or equal to 9
in three rolls of a die ?

Answer by ewatrrr(24785)


Hi
three rolls of a die:  6^3 = 216 ways they can be rolled.
•	3 = 1 + 1 + 1 , 1 way
•	4 = 1 + 1 + 2 , 3 ways * 3!/2!
•	5 = 1 + 1 + 3 = 2 + 2 + 1,  6 ways 3!/2! + 3!/2!
•	6 = 1 + 1 + 4 = 1 + 2 + 3 = 2 + 2 + 2  10ways  3!/2! + 3! + 1
•	7 = 1 + 1 + 5 = 2 + 2 + 3 = 3 + 3 + 1 = 1 + 2 + 4 , 15  ways etc
•	8 = 1 + 1 + 6 = 2 + 3 + 3 = 4 + 3 + 1 = 1 + 2 + 5 = 2 + 2 + 4, 21 ways
•	9 = 6 + 2 + 1 = 4 + 3 + 2 = 3 + 3 + 3 = 2 + 2 + 5 = 1 + 3 + 5 = 1 + 4 + 4, 25 ways
P(sum ≤ 9) = 1+3 +6+10+15+21+25 = 81/6^3 = 81/216 = 3/8 
Wish You the Best in your Studies.